Two types of milk, one that has 1.5% butterfat, and the other that has 4.5% of butterfat, are mixed. How many liters of these two different kinds of milk are to be mixed together to produce 100 liters of low-fat milk, which has 2.5% butterfat?
Answer by richwmiller(17219)

3.25% Whole milk or Regular Milk
about 2% =2% or Reduced fat
about 1% =1% or Low fat
0.0 – 0.5% =Skim milk or nofat
http://en.wikipedia.org/wiki/Fat_content_of_milk
Where did you get 4.5 % butterfat milk? Normal is 3.25%
Low fat milk is 1%
reduced is 2%
.015a+.045b=.025*100,
a+b=100
a= 66 2/3 b=33 1/3